# Multi-LLM Thematic Analysis with Dual Reliability Metrics: Combining Cohen’s Kappa and Semantic Similarity for Qualitative Research Validation

### v1

@@ 1 Introduction
Recent LLM-based systems such as QualIT (bhaduri2024qualitative, ) focus on topic modeling and key-phrase extraction, achieving 70% topic coherence on benchmark datasets. However, these approaches differ fundamentally from comprehensive thematic analysis as defined by Braun and Clarke (braun2006using, ), which involves iterative interpretation, contextualization, and affective understanding. Comparative studies (bennis2025advancing, ; sakaguchi2025evaluating, ) evaluating nine generative models reveal significant performance variation across models and highlight cultural interpretation challenges, particularly in non-Western contexts. These findings underscore a critical gap: existing LLM approaches lack systematic validation mechanisms for reliability assessment.
@@ 2 Related Work
Traditional Reliability Assessment. Qualitative research relies on inter-rater reliability to establish trustworthiness (braun2006using, ). Cohen’s kappa measures agreement between two coders, with values interpreted as follows: $\kappa<0.40$ (poor), 0.40-0.60 (moderate), 0.60-0.80 (substantial), $\kappa>0.80$ (excellent). However, kappa requires exact categorical matches and cannot capture semantic equivalence. Studies report that even trained coders often achieve only moderate agreement, necessitating extensive discussion to resolve discrepancies (citation).
@@
LLM-Based Qualitative Analysis. Recent work explores LLM applications in qualitative research. QualIT (bhaduri2024qualitative, ) integrates LLMs with clustering for topic modeling, extracting key phrases and performing hierarchical clustering to achieve 70% topic coherence on benchmark datasets. However, this approach focuses on topic extraction rather than comprehensive thematic analysis.
@@
Alternative frameworks propose human-LLM collaboration models. Rana and Asad (rana2023llm, ) introduce ”LLM-in-the-loop,” using in-context learning with GPT-3.5 to reduce labor requirements while maintaining human oversight. Schlagwein (schlagwein2024role, ) proposes four conversational roles (managers, teachers, colleagues, advocates) for researchers working with LLM chatbots, emphasizing reflexive practice. Landers and Behrend (landers2024reconciling, ) employ Retrieval-Augmented Generation (RAG)-based approaches for interview transcript analysis, focusing on methodological rigor. The A Human-AI Collaborative Thematic Analysis framework using Multi-Agent (TAMA) framework (alhakeem2025tama, ) applies multi-agent LLMs specifically to clinical interviews, demonstrating domain-specific adaptations.
@@
Lindh and Messina (lindh2023performing, ) demonstrate that LLMs can infer main themes but highlight limitations in capturing latent interpretations—themes requiring deep contextual understanding. Turobov et al. (turobov2024using, ) evaluate ChatGPT for thematic analysis, finding promise but recommending human oversight. Fulgencio (fulgencio2024ai, ) identifies benefits while noting cultural context limitations.
@@
Comparative LLM Studies. Bennis and Mouwafaq (bennis2025advancing, ) conduct a comparative study of nine generative models on medical data, revealing significant performance variation across models (something more specific). Sakaguchi et al. (sakaguchi2025evaluating, ) compare ChatGPT with human researchers in Japanese clinical contexts, highlighting cultural interpretation challenges that LLMs struggle to address. Zhang et al.’s LLM-Assisted Thematic Analysis (LATA) study (zhang2025lata, ) compares GPT-4 and Gemini outputs with manually analyzed outcomes, achieving cosine similarity scores up to 0.76—validating semantic similarity as a viable metric but also revealing model-dependent variation. Gupta et al. (gupta2025utilizing, ) examine LLMs for focus group transcript analysis, demonstrating potential but noting reliability concerns.
@@
Prompt Engineering and Quality. Prompt design critically impacts analysis quality. Braun and Clarke (braunClarke2023prompt, ) provide reproducible prompt engineering approaches aligned with their five-phase framework, with empirical evaluation against established quality criteria. Sanford et al. (sanford2024enhancing, ) systematically evaluate prompt engineering techniques using locally hosted Llama 3.1 models, demonstrating that structured prompts significantly improve thematic coherence. Nelson (nelson2025reflecting, ) tests offline LLMs through reflexive thematic analysis phases, identifying limitations of base models and proposing prompt strategies for improvement.
@@
Validation Metrics. Novel validity metrics emerge for LLM-assisted analysis. Patel et al. (patel2024reflections, ) propose initial thematic saturation (ITS) as a validity metric, measuring when LLMs reach analytical saturation in initial coding. Chen et al. (chen2025codebook, ) examine codebook reduction and saturation patterns, providing insights into how LLMs handle iterative coding processes.
@@
Systematic Evidence. Kumar et al. (kumar2025large, ) provide a comprehensive systematic mapping of LLM applications in qualitative research across diverse fields, application contexts, and evaluation metrics. Their review reveals heterogeneous approaches with limited standardization, underscoring the need for systematic validation frameworks.
@@
These studies reveal a critical gap: existing LLM approaches lack systematic validation mechanisms. Single-run analyses provide no reliability indicators, and multi-model studies focus on performance comparison rather than developing validation frameworks. Our work addresses this gap through ensemble validation with quantified reliability metrics, building on the semantic similarity validation demonstrated by Zhang et al. (zhang2025lata, ) while extending it through multi-run consensus.
@@ 3.1 Ensemble Validation Framework
Our framework conducts six independent analytical runs with fixed random seeds (42, 123, 456, 789, 1011, 1213), analogous to K-fold cross-validation in machine learning. This design choice is grounded in statistical theory and practical considerations.
@@ 5.2 Comparison with Existing Approaches
Our dual-metric validation extends recent work: Zhang et al.’s LATA (zhang2025lata, ) achieved 0.76 cosine similarity between LLM and human analyses; our inter-run consistency exceeds this (0.92-0.95), suggesting ensemble methods may achieve higher reliability than single-run human-AI comparison. The TAMA framework (alhakeem2025tama, ) employs multi-agent architectures; our single-agent multi-run approach provides simpler implementation with comparable reliability. QualIT (bhaduri2024qualitative, ) focuses on key-phrase extraction; our structure-agnostic design supports full thematic analysis with arbitrary output formats.
@@ 5.3 Interpretation Guidelines
Cohen’s Kappa. Following Landis-Koch criteria: $\kappa>0.80$ (almost perfect), 0.60-0.80 (substantial), 0.40-0.60 (moderate). Our results ($\kappa=0.84$-$0.91$) achieve ”almost perfect” reliability across all three LLMs, validating the ensemble approach for rigorous qualitative research.
Cosine Similarity. Interpret as percentage semantic overlap: >>90% (high consistency), 80-90% (moderate), <<80% (low, warrants review). Our results (92-95%) demonstrate strong convergence. Kappa range (spread between min/max pairwise kappa) indicates stability: <<0.25 (stable), 0.25-0.40 (moderate variation), >>0.40 (high variation requiring investigation).
Consensus Thresholds. Default 50% (3/6 runs) balances sensitivity and specificity. Adjust based on context: 67% (4/6) for conservative high-stakes research, 33% (2/6) for exploratory analysis. High-confidence themes ($\geq$83%, 5-6/6 runs) require minimal human review; moderate-confidence themes (50-66%, 3-4/6 runs) warrant researcher judgment.
@@ 5.4 Limitations
Cultural and Domain Boundaries. LLMs encode training data biases (sakaguchi2025evaluating, ). Our framework aids bias detection (biased themes appearing in 1-2/6 runs flag for review), but cannot eliminate it. Performance on non-English, non-Western, or highly specialized domain data requires systematic evaluation.
@@ 5.5 Future Work
Human-AI Comparison. Comparison against human coders on identical datasets, measuring kappa agreement between AI consensus themes and human-coded themes. Zhang et al. (zhang2025lata, ) achieved 0.76 similarity; our inter-run consistency (0.92-0.95) suggests potential for high human-AI agreement.
Adaptive Run Configuration. Implementing thematic saturation metrics (patel2024reflections, ; chen2025codebook, ) to determine optimal run counts dynamically. If new themes cease emerging after N runs, stop analysis rather than using fixed N=6.
@@ Acknowledgments
This research was conducted at Aza Lab at Department of Psychiatry at Yale University and the Center of Collective Healing at Howard University. Funding provided by internal grants to the AZA Lab. The complete source code is openly available on GitHub at https://github.com/NileshArnaiya/LLM-Thematic-Analysis-Tool, with live deployment at https://azalab-llm-tool.vercel.app/. The ketamine art therapy interview transcript used for evaluation is available in the repository under /transcript.txt. We thank the participating therapist for sharing insights on integrating expressive arts with ketamine-assisted psychotherapy.
AA designed and supervised project, wrote and edited mansucript. NA developed model, wrote and implemented code, and drafted manuscript. SA collected and provided qualitative data

### v2

@@ 1 Introduction
Recent LLM-based systems have evolved from simple prompting to complex architectures. Frameworks like QualIT (kapoor2024qualit, ) utilize iterative clustering, while multi-agent systems like Thematic-LM (qiao2025thematic, ) and TAMA (xu2025tama, ) leverage agentic interactions to enhance the quality of thematic extraction.
However, a critical gap remains in reliability quantification. While these methods employ multiple inference steps to refine outputs, they typically present the final result as a definitive conclusion without assessing its stochastic stability. They rarely quantify how sensitive the analysis is to the model’s inherent probabilistic nature. Without measuring variance across independent runs, it remains unclear whether a generated theme is a robust finding or an artifact of a specific interaction path.
As noted in reasoning tasks (wang2022self, ), single-point generation fails to capture the model’s full interpretive distribution. Consequently, single-run analyses lack statistical robustness and fail to distinguish between stable thematic signals and hallucinatory noise.
@@ 2 Related Work
Traditional Reliability Assessment. Qualitative research relies on inter-rater reliability to establish trustworthiness (braun2006using, ). Cohen’s kappa measures agreement between two coders, with values interpreted as follows: $\kappa<0.40$ (poor), 0.40-0.60 (moderate), 0.60-0.80 (substantial), $\kappa>0.80$ (excellent). However, kappa requires exact categorical matches and cannot capture semantic equivalence. Studies report that even trained coders often achieve only moderate agreement, necessitating extensive discussion to resolve discrepancies (shrout1979intraclass, ).
@@
LLM-Based Qualitative Analysis. Recent work explores LLM applications in qualitative research. QualIT (kapoor2024qualit, ) integrates LLMs with clustering for topic modeling, extracting key phrases and performing hierarchical clustering to achieve 70% topic coherence on benchmark datasets. However, this approach focuses on topic extraction rather than comprehensive thematic analysis.
@@
Alternative frameworks propose human-LLM collaboration models. Dai et al. (dai2023llm, ) introduce ”LLM-in-the-loop,” using in-context learning with GPT-3.5 to reduce labor requirements while maintaining human oversight. Thominet (thominet2024role, ) proposes four conversational roles (managers, teachers, colleagues, advocates) for researchers working with LLM chatbots, emphasizing reflexive practice. Bhaduri et al. (bhaduri2024reconciling, ) employ Retrieval-Augmented Generation (RAG)-based approaches for interview transcript analysis, focusing on methodological rigor. The A Human-AI Collaborative Thematic Analysis framework (TAMA) (xu2025tama, ) applies multi-agent LLMs specifically to clinical interviews. Similarly, Thematic-LM (qiao2025thematic, ) employs a multi-agent system to scale thematic analysis, utilizing agent interaction to enhance interpretative depth.
@@
De Paoli (depaoli2024performing, ) demonstrates that LLMs can infer main themes but highlight limitations in capturing latent interpretations—themes requiring deep contextual understanding. Turobov et al. (turobov2024using, ) evaluate ChatGPT for thematic analysis, finding promise but recommending human oversight. Christou (christou2024thematic, ) identifies benefits while noting cultural context limitations.
@@
Comparative LLM Studies. Bennis and Mouwafaq (bennis2025advancing, ) conduct a comparative study of nine generative models on Cutaneous Leishmaniasis medical data, revealing significant performance variation across models in thematic coherence and clinical relevance. Sakaguchi et al. (sakaguchi2025evaluating, ) compare ChatGPT with human researchers in Japanese clinical contexts, highlighting cultural interpretation challenges that LLMs struggle to address. Wang et al.’s LATA study (wang2025lata, ) compares GPT-4 and Gemini outputs with manually analyzed outcomes, achieving cosine similarity scores up to 0.76—validating semantic similarity as a viable metric but also revealing model-dependent variation. Wong et al. (wong2025utilizing, ) examine LLMs for focus group transcript analysis, demonstrating potential but noting reliability concerns.
@@
Prompt Engineering and Quality. Prompt design critically impacts analysis quality. Martinez Montes et al. (montes2025largelanguagemodelsthematic, ) provide reproducible prompt engineering approaches aligned with their five-phase framework, with empirical evaluation against established quality criteria. Meyer et al. (meyer2025enhancing, ) systematically evaluate prompt engineering techniques using locally hosted Llama 3.1 models, demonstrating that structured prompts significantly improve thematic coherence. Vikan et al. (vikan2025reflecting, ) tests offline LLMs through reflexive thematic analysis phases, identifying limitations of base models and proposing prompt strategies for improvement.
@@
Validation Metrics. Novel validity metrics emerge for LLM-assisted analysis. De Paoli and Mathis (depaoli2025reflections, ) propose initial thematic saturation (ITS) as a validity metric, measuring when LLMs reach analytical saturation in initial coding. De Paoli and Mathis (depaoli2025codebook, ) examine codebook reduction and saturation patterns, providing insights into how LLMs handle iterative coding processes.
@@
Systematic Evidence. Barros et al. (barros2024large, ) provide a comprehensive systematic mapping of LLM applications in qualitative research across diverse fields, application contexts, and evaluation metrics. Their review reveals heterogeneous approaches with limited standardization, underscoring the need for systematic validation frameworks.
@@
These studies reveal a critical gap: existing LLM approaches lack systematic validation mechanisms. Single-run analyses provide no reliability indicators, and multi-model studies focus on performance comparison rather than developing validation frameworks. Our work addresses this gap through ensemble validation with quantified reliability metrics, building on the semantic similarity validation demonstrated by Zhang et al. (wang2025lata, ) while extending it through multi-run consensus.
@@ 3.1 Ensemble Validation Framework
Our framework employs a semantic Monte Carlo simulation approach utilizing six independent analytical runs with fixed random seeds (42, 123, 456, 789, 1011, 1213). Rather than viewing these as mere repetitions, we treat each run as an independent sample from the model’s posterior distribution over potential thematic interpretations. This design choice is grounded in statistical theory and practical considerations.
@@ 4.5 Comparison with Existing Frameworks
Traditional cost includes synthesis labor across all documents; AI framework costs reflect per-document analysis without cross-document synthesis.
@@ 5.2 Comparison with Existing Approaches
Our dual-metric validation extends recent work in two key directions: internal consistency and methodological parsimony.
Comparison with Single-Run Approaches. Wang et al.’s LATA (wang2025lata, ) achieved 0.76 cosine similarity between LLM and human analyses. In contrast, our inter-run consistency (0.92–0.95) demonstrates strong internal reliability—serving as a proof of concept for consistency across multiple runs. However, high internal consistency does not automatically guarantee external validity (accuracy against human ground-truth). Without human baseline comparison on this specific dataset, we cannot claim our ensemble achieves higher human-AI agreement than single-run approaches; rather, we establish that ensemble methods provide quantifiable, reproducible consistency metrics.
Comparison with Multi-Agent Systems (Thematic-LM & TAMA). Regarding architectural differences, frameworks like Thematic-LM (qiao2025thematic, ) and TAMA (xu2025tama, ) employ multi-agent architectures to simulate human-like discourse. It is important to distinguish our goal from theirs:
  • Performance vs. Validation: Multi-agent approaches utilize diversity (e.g., debating agents) primarily to optimize performance—aiming to generate deeper themes. In contrast, our ensemble framework utilizes diversity (via seeds) to validate reliability. We do not seek to artificially enhance the themes through debate, but rather to rigorously test their reproducibility under controlled stochastic conditions.
  • Methodological Parsimony: While effective, multi-agent systems introduce significant architectural complexity and opacity. We argue for parsimony. By adopting a single-agent ensemble approach inspired by self-consistency decoding (wang2022self, ), we achieve high reliability ($\kappa>0.90$) through independent validation rather than complex inter-agent dynamics. This explicitly isolates the model’s internal consistency from interaction artifacts.
@@ 5.3 Interpretation Guidelines
Cohen’s Kappa. Following Landis-Koch criteria: $\kappa>0.80$ (almost perfect), 0.60–0.80 (substantial), 0.40–0.60 (moderate). Our results ($\kappa=0.84$$0.91$) achieve ”almost perfect” reliability across all three LLMs, validating the ensemble approach for rigorous qualitative research.
Cosine Similarity. Interpret as percentage semantic overlap: $90\%$ (high consistency), 80–90% (moderate), $80\%$ (low, warrants review). Our results (92–95%) demonstrate strong convergence. Kappa range (spread between min/max pairwise kappa) indicates stability: <0.25<0.25 (stable), 0.25–0.40 (moderate variation), >0.40>0.40 (high variation requiring investigation).
Consensus Thresholds. Default 50% (3/6 runs) balances sensitivity and specificity. Adjust based on context: 67% (4/6) for conservative high-stakes research, 33% (2/6) for exploratory analysis. High-confidence themes ($\geq$83%, 5–6/6 runs) require minimal human review; moderate-confidence themes (50–66%, 3–4/6 runs) warrant researcher judgment.
@@ 5.4 Limitations
Cultural and Domain Boundaries. LLMs encode training data biases (sakaguchi2025evaluating, ). Our framework aids bias detection (biased themes appearing in 1–2/6 runs flag for review), but cannot eliminate it. Performance on non-English, non-Western, or highly specialized domain data requires systematic evaluation.
@@ 5.5 Future Work
Human-AI Comparison. Comparison against human coders on identical datasets, measuring kappa agreement between AI consensus themes and human-coded themes. Wang et al. (wang2025lata, ) achieved 0.76 similarity; our inter-run consistency (0.92-0.95) suggests potential for high human-AI agreement.
Adaptive Run Configuration. Implementing thematic saturation metrics (depaoli2025reflections, ; depaoli2025codebook, ) to determine optimal run counts dynamically. If new themes cease emerging after N runs, stop analysis rather than using fixed N=6.
@@ Acknowledgments
This research was conducted at Aza Lab at Department of Psychiatry at Yale University and the Center of Collective Healing at Howard University. Funding provided by internal grants to the AZA Lab. The complete source code is openly available on GitHub at https://github.com/NileshArnaiya/LLM-Thematic-Analysis-Tool, with live deployment at https://azalab-llm-tool.vercel.app/. The ketamine art therapy interview transcript used for evaluation is available in the repository under /transcript.txt. We thank the participating therapist for sharing insights on integrating expressive arts with ketamine-assisted psychotherapy. We are grateful to Leor Roseman for his valuable guidance on model usability and implementation. We thank Seyi Adeyinka for collecting and providing the qualitative dataset used in this study. We also thank Marik Hazan for their fruitful discussions.
Aza Allsop designed and supervised the project, wrote and edited the manuscript. Nilesh Jain developed the model, wrote and implemented the code, and drafted the manuscript. Hyungil Suh contributed to the methodological conceptualization, refined the validation framework arguments, and performed critical review. Leor Roseman helped refine and guide model usability and implementation.